OpenStudy (anonymous):

x^2+y^2=25 36x^2-4y^2=144

OpenStudy (lalaly):

multiply the first equation by 4 4x^2+4y^2=100 36x^2-4y^2=144 subtract both equations -32x^2=-44 x^2=8/11 substitute in first equation 8/11 +y^2=25 y^2= 25-8/11 y^2=267/11